OpenStudy (anonymous):

multiply (3√(5+1)(2√(5-3)

jhonyy9 (jhonyy9):

30-9sqrt5+2sqrt5-3=27-7sqrt5

OpenStudy (anonymous):

ans is 12sqr(3)

jhonyy9 (jhonyy9):

so if there are 3sqrt(5+1)*2sqrt(5-3) so than this is equal 3sqrt6 * 2sqrt2 =6sqrt12 =6sqrt(4*3)=6*2sqrt3 =12sqrt3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yea.....

jhonyy9 (jhonyy9):

ok good luck bye